OpenStudy (anonymous):

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

Parth (parthkohli):

Use the Pythagorean theorem. \( \color{Black}{\Rightarrow 1^2 + 1^2 = c^2 }\)

OpenStudy (anonymous):

Isosceles means that In a triangle any two sides are equal in length.. As it is a right angled triangle, so its two sides are of 1 unit length and there is one hypotenuse length of which is to be found.. So,\[H^2 = 1^2 + 1^2\] \[H = \sqrt{2}\] Hypotenuse Length = \[\sqrt{2}\] units..